# Large Language Models on Fine-grained Emotion Detection Dataset with Data Augmentation and Transfer Learning

### 10 Mar 2024 v1

@@ 6 Conclusion and Future Work
Data augmentation, particularly when applied to under-performing categories, has proven to be a potent tool in balancing the dataset Dong et al. (2021) and enhancing model accuracy. Among the augmentation techniques, PROT augmented dataset has emerged as the superior strategy, outperforming other methods in improving the classification results of our models.
On top of that, our study also highlights the complexity and sensitivity of machine learning tasks to their respective datasets. The initially hypothesized superiority of the RoBERTa model was not realized, underscoring the need for task-specific model selection and configuration. Furthermore, we observed that while data augmentation can indeed propel performance forward, it requires a nuanced application to prevent exacerbating existing data imbalances.

### 09 Apr 2024 v2

@@ 6 Conclusion and Future Work
Data augmentation, particularly when applied to under-performing categories, has proven to be a potent tool in balancing the dataset Dong et al., (2021) and enhancing model accuracy. Among the augmentation techniques, PROT augmented dataset has emerged as the superior strategy, outperforming other methods in improving the classification results of our models.
On top of that, our study also highlights the complexity and sensitivity of machine learning tasks to their respective datasets. The initially hypothesized superiority of the RoBERTa model was not realized, underscoring the need for task-specific model selection and configuration. Furthermore, we observed that while data augmentation can indeed propel performance forward, it requires a nuanced application to prevent exacerbating existing data imbalances. Also, we identified three primary issues contributing to GPT-4’s inferior performance metrics: hallucination, over-labelling, and over-interpretation when applying GPT-4 to GoEmotions dataset.